# Large Language Models as Virtual Survey Respondents: Evaluating Sociodemographic Response Generation

### v1

@@ 5. Discussions
Our experiments reveal critical insights into optimizing LLMs for survey simulation across three dimensions: reasoning-oriented distilled models, contextual modulation, and in-context learning dynamics. These findings provide actionable guidelines for deploying LLMs in sociological research while highlighting remaining challenges.
@@ 5.1. Leveraging Reasoning-Oriented Distilled Models
As survey simulation requires nontrivial reasoning over attributes, we investigate distilled LLMs explicitly optimized for reasoning, such as those from the DeepSeek-R1-Distill family (Guo et al., 2025). To evaluate their effectiveness, we compare sociodemographic response generation in the PAS scenario on the YPS dataset between standard LLaMA models (3.1-8B and 3.3-70B) and their reasoning-optimized distilled counterparts (DeepSeek-R1-Distill-Llama-8B and -70B). Figure 5 reveals that the reasoning-oriented distillation leads to a notable performance improvement for the 8B LLaMA model, yielding a relative increase of 40.6% and indicating enhanced response generation capabilities. However, for the 70B LLaMA model, the distillation does not provide a significant advantage, instead resulting in a relative decrease of 22.9%.
These findings suggest that while the reasoning-oriented distillation effectively boosts the performance of smaller models, its benefits diminish in larger models under the PAS setting. One possible explanation is that larger models, with their extensive parameter space, already possess the capacity to generate high-quality responses without requiring distillation (Kaplan et al., 2020), and applying further reasoning-oriented distillation may induce excessive internal deliberation in complex tasks, potentially impairing output accuracy or coherence (Cuadron et al., 2025). In contrast, smaller models, limited by fewer parameters, benefit more from the distillation as it helps streamline their processing and enhances their ability to simulate accurate sociodemographic responses (Ramesh et al., 2025).
@@ 5.2. Context-Aware Prediction Enhancement
To demonstrate the impact of integrating prompt-based contextual information, we show that embedding precise contextual cues, such as explicit references to domains, timeframes, or real-world policy frameworks, significantly enhances prediction accuracy in sociodemographic response generation. Our experiments with the LLaMA 3.1-8B model reveal that the contextual information enables the model to activate latent knowledge and produce predictions that are more closely aligned with empirical scenarios. To evaluate the effectiveness of context-aware prediction, we concentrated on four key domains of public opinion concerning government spending in the GSS dataset. Contextual information was introduced to guide the model’s response generation. For example, the prompt “Considering the environmental policies of the United States in 2022” was used to provide contextual grounding in the environmental domain.
Incorporating contextual prompts led to improved prediction performance across the four domains, namely environmental protection (natenvir), big city issues (natcity), welfare programs (natfare), and social security (natsoc), with relative gains ranging from 6.67% to 23.62% (see Figure 6). These results establish context integration as a critical requirement for survey simulation systems, governed by two synergistic mechanisms: first, contextual anchoring overrides LLMs’ tendency for shallow pattern matching when addressing complex societal issues; second, domain-specific knowledge activation closes the abstraction gap between model outputs and tangible policy dynamics (Sahoo et al., 2024). Crucially, these findings mandate a paradigm shift reframing context-aware scaffolding not as an optional add-on, but as a necessary component to be systematically implemented in high-stakes applications. This shift ensures that LLMs evolve from mere statistical pattern extractors into contextually grounded reasoning systems capable of supporting societal decision modeling.
@@ 5.3. Mitigating In-Context Learning Bias
We manipulated example distributions using three configurations and analyzed their output distributions, highlighting the ratio of option 2: (i) all option 2 (82.38%), (ii) no option 2 (57.76%), and (iii) random selection (80.01%). Notably, configuration (ii) caused a 24.62 percentage point decline compared to (i) (see Figure 7), underscoring the strong influence of embedded input distributions on model outputs. Statistical evidence from our experiments reveals a strong dependency between example balance and predictive robustness in our few-shot evaluation, directly motivating our adoption of strict example balancing measures. Crucially, these findings ground our operational decision to enforce balanced few-shot demonstrations in standardized prompts, ensuring counterfeit-resistant predictions aligned with real-world policy dynamics.
@@ 6. Conclusion
To evaluate large language models in sociodemographic response generation, we propose two paradigms: Partial Attribute Simulation (PAS) and Full Attribute Simulation (FAS), each targeting distinct levels of inference and data availability. PAS addresses individual-level attribute inference under incomplete profiles, while FAS targets population-level synthesis from limited demographic signals, supporting analyses in data-sparse settings. On this basis, we further introduce LLM-S3 (Large Language Model-based Sociodemographic Survey Simulation), a comprehensive benchmark suite spanning eleven datasets across four domains. Our evaluation reveals systematic differences across model series and prompting strategies, establishing a foundation for rigorous LLM assessment and advancing trustworthy survey simulations in computational social science.
@@ A.1. Details of Survey Datasets from Different Domains
[…]
@@ A.3. Attribute Configuration & Detailed Performance
[…]

### v2

@@ 5. Discussions
The baseline experiments establish the capabilities and failure modes of LLMs for survey simulation tasks. In this section, we investigate three optimization dimensions that directly affect simulation fidelity: (1) the role of reasoning-oriented model distillation in enhancing attribute inference, (2) the impact of contextual prompting on improving response accuracy, and (3) the sensitivity of few-shot evaluation to example balance. Together, these analyses provide actionable guidelines for deploying LLMs in sociological research while surfacing remaining challenges.
It is important to emphasize that our evaluation focuses on statistical and distributional similarity rather than behavioral fidelity or construct validity. This distinction has direct implications for how the findings below should be interpreted: improvements in accuracy or KL-based scores reflect better alignment with empirical distributions, not necessarily a deeper understanding of the social processes that generate those distributions.
@@ 5.1. Leveraging Reasoning-Oriented Distilled Models
Inferring survey attributes from incomplete respondent profiles requires nontrivial reasoning over attribute interdependencies. This section investigates whether distilled LLMs explicitly optimized for reasoning, such as those from the DeepSeek-R1-Distill family (Guo et al., 2025), offer advantages over standard instruction-tuned models in survey simulation settings. To evaluate their effectiveness, we compare sociodemographic response generation on the YPS dataset between standard LLaMA models (3.1-8B and 3.3-70B) and their reasoning-optimized distilled counterparts (DeepSeek-R1-Distill-Llama-8B and -70B). Figure 7 reveals that the reasoning-oriented distillation leads to a notable performance improvement for the 8B LLaMA model, yielding a relative increase of 40.6% and indicating enhanced response generation capabilities. However, for the 70B LLaMA model, the distillation does not provide a significant advantage, instead resulting in a relative decrease of 22.9%.
These findings suggest that while the reasoning-oriented distillation effectively boosts the performance of smaller models, its benefits diminish in larger models. One possible explanation is that larger models, with their extensive parameter space, already possess the capacity to generate high-quality responses without requiring distillation (Kaplan et al., 2020), and applying further reasoning-oriented distillation may induce excessive internal deliberation in complex tasks, potentially impairing output accuracy or coherence (Cuadron et al., 2025). In contrast, smaller models, limited by fewer parameters, benefit more from the distillation as it helps streamline their processing and enhances their ability to simulate accurate sociodemographic responses (Ramesh et al., 2025). These results suggest that practitioners working with resource-constrained deployments can achieve meaningful accuracy gains by substituting standard small models with their reasoning-distilled counterparts, whereas large-model deployments should be evaluated on a case-by-case basis before adopting distillation.
@@ 5.2. Context-Aware Prediction Enhancement
Survey responses are inherently situated within specific sociopolitical, temporal, and cultural contexts. A respondent’s attitudes toward environmental policy, for example, cannot be fully understood without reference to the prevailing policy landscape and recent events. This section investigates whether explicitly contextualizing prompts with domain-relevant information improves the fidelity of LLM-generated survey responses. To demonstrate the impact of integrating prompt-based contextual information, we show that embedding precise contextual cues, such as explicit references to domains, timeframes, or real-world policy frameworks, significantly enhances prediction accuracy in sociodemographic response generation. Our experiments with the LLaMA 3.1-8B model reveal that the contextual information enables the model to activate latent knowledge and produce predictions that are more closely aligned with empirical scenarios. To evaluate the effectiveness of context-aware prediction, we concentrated on four key domains of public opinion concerning government spending in the GSS dataset. Contextual information was introduced to guide the model’s response generation. For example, the prompt “Considering the environmental policies of the United States in 2022” was used to provide contextual grounding in the environmental domain.
Incorporating contextual prompts led to improved prediction performance across the four domains, namely environmental protection (natenvir), big city issues (natcity), welfare programs (natfare), and social security (natsoc), with relative gains ranging from 6.67% to 23.62% (see Figure 8). These results establish context integration as a critical requirement for survey simulation systems, governed by two synergistic mechanisms: first, contextual anchoring overrides LLMs’ tendency for shallow pattern matching when addressing complex societal issues; second, domain-specific knowledge activation closes the abstraction gap between model outputs and tangible policy dynamics (Sahoo et al., 2024). These findings suggest that survey simulation practitioners should embed domain-specific contextual cues in prompts whenever the target survey domain is known. The consistent performance gains observed across diverse policy domains indicate that context-aware prompting is a broadly applicable strategy for improving simulation fidelity, regardless of the specific task setting. This shift ensures that LLMs evolve from mere statistical pattern extractors into contextually grounded reasoning systems capable of supporting societal decision modeling.
@@ 5.3. Mitigating In-Context Learning Bias
Few-shot learning enables LLMs to adapt to specific tasks with limited examples, but the selection and balance of these examples can systematically bias model outputs. In survey simulation, where demographic representation and response distributions carry substantive meaning, such biases pose significant validity threats. This section examines how example distribution affects prediction accuracy in few-shot survey simulation settings.
@@
We manipulated example distributions using three configurations and analyzed their output distributions, highlighting the ratio of option 2: (i) all option 2 (82.38%), (ii) no option 2 (57.76%), and (iii) random selection (80.01%). Notably, configuration (ii) caused a 24.62 percentage point decline compared to (i) (see Figure 9), underscoring the strong influence of embedded input distributions on model outputs. Statistical evidence from our experiments reveals a strong dependency between example balance and predictive robustness in our few-shot evaluation, directly motivating our adoption of strict example balancing measures. These findings suggest that practitioners employing few-shot learning for survey simulation should draw demonstrations from balanced class distributions and treat the sensitivity of model outputs to example composition as a systematic evaluation criterion rather than an implementation detail.
5.4. Ethical Considerations
The use of LLMs to generate synthetic survey responses raises four ethical considerations that must be addressed before these technologies can be widely adopted in sociological research:
(1) Bias amplification. Our experiments demonstrate that LLMs can amplify existing biases present in training data. For example, predictions about income based on education level may exhibit more pronounced gender or racial disparities than those observed in ground-truth data. This risk necessitates rigorous bias audits before deploying LLM-based survey simulation tools.
(2) Representation and generalizability. The datasets evaluated in this study are predominantly drawn from North American and European contexts, reflecting the current availability of large-scale public survey data rather than a fundamental constraint of the evaluation framework. The benchmark is designed to be extensible: as cross-cultural and multilingual survey datasets become available, the same evaluation protocol can be applied to assess LLM performance in diverse sociological contexts.
(3) Risk of misuse or misrepresentation. Without proper disclosure, LLM-generated responses could be mistaken for genuine human survey data, potentially leading to flawed research conclusions or policy decisions. Practitioners should explicitly disclose when synthetic data has been used and clarify the limitations of LLM-generated responses.
(4) Transparency and reproducibility. Synthetic data generation processes should be transparent and reproducible. In contexts with limited data availability, synthetic data should be used as a complement to, rather than a replacement for, genuine survey data.
5.5. Limitations and Future Research Directions
Despite the strengths of our framework and benchmark, our work has several limitations that suggest directions for future research: (1) Our evaluation focuses on statistical and distributional similarity rather than behavioral fidelity or construct validity, which are important considerations in social science research; (2) The datasets evaluated in this study are predominantly drawn from North American and European contexts; extending the benchmark to cross-cultural and multilingual settings is an important direction for future work; (3) While we have examined partial and full attribute simulation as complementary approaches, further research is needed to explore hybrid methods that combine elements of both; (4) Our analysis of failure cases, while significant, is not exhaustive, and future work should continue to identify and address the limitations of LLM-based survey simulation.
@@ 6. Conclusion
Questionnaire-based surveys remain indispensable for understanding public sentiment and informing evidence-based policy, yet their high cost, lengthy timelines, and declining response rates motivate the exploration of LLM-based simulation as a complementary analytical tool. To support systematic evaluation of this emerging paradigm, this work formalizes two task settings that reflect distinct real-world data availability conditions: Partial Attribute Simulation (PAS), where individual demographic records are available but attitudinal attributes are missing, and Full Attribute Simulation (FAS), where only aggregate population statistics are available and a synthetic population must be generated from scratch. A unified benchmark, LLM-S3, spanning eleven real-world datasets across four sociological domains, is curated to enable consistent, reproducible evaluation across models and settings.
Experiments across four representative LLMs yield several actionable insights. In PAS, the benchmark reveals which survey constructs maintain detectable statistical associations with available demographic profiles, enabling practitioners to identify suitable simulation targets before deployment. In FAS, the evaluation surfaces how model scale and alignment strategy jointly shape the fidelity of synthetic population generation. Taken together, these findings position LLM-S3 as a diagnostic instrument that helps researchers calibrate expectations, select appropriate models, and design more effective prompting strategies for LLM-assisted survey research. We hope this benchmark serves as a foundation for future work on broadening the sociological scope, improving cross-cultural coverage, and advancing the methodological rigor of LLM-based survey simulation.
@@ A.1. Details of Survey Datasets from Different Domains
| Domain | Survey Name | Abbr. | Year | Description | # Cases | # Attrs. |
@@ A.3. Attribute Configuration & Detailed Performance
| Dataset | Simulation Objectives | Input Attributes | Target Attributes |